1. (20 point compute the ionic (a) 0.02m NaC (b) 0.01m K2SOA(ag)+ 0.03m KBr(a) ag)+ 0.03m NaOH(a)
Answer #1

A) ionic strength= 0.5 summation CiZi^2

= 0.5*(0.02+0.03+0.02+0.03) = 0.05

B) ionic strength= 0.5*(0.02+0.03+0.03+0.01*2^2)

= 0.06